# Feynman Integral Reduction without Integration-By-Parts

### v1

@@ 1 Introduction
The paper is organized as follows. In Section 2, we introduce the equivalence relations of integral contours in Feynman parameterization and use several simple examples to demonstrate our reduction method. In Section 3, we present our general method for one-loop integral reduction, and provide recursive formulas that can be easily implemented in computer algebra. In Section 4, we demonstrate our method through several examples. In Section 5, we provide a summary and explore the potential extension of this method to higher loops, discussing the new challenges that may arise in future applications.
@@ 2.1 Feynman parametrization and equivalence classes of integration domains
where $\mathcal{X}$ and $\mathcal{Y}$ are non-negative homogeneous functions of degree 00 and $1$ in the variables $\{\alpha_{j}\}$, and we also require that the equation $\mathcal{X}-\mathcal{Y}=0$ has only one solution for any variable $\alpha_{j}$ in the whole integration domain. The original representation (2) then corresponds to the specific choice $\mathcal{X}=1$ and $\mathcal{Y}=\sum_{j\in\mathcal{S}}\alpha_{j}$. Note that by choosing $\mathcal{X}=1$ and $\mathcal{Y}=\alpha_{0}$, we can integrate out $\alpha_{0}$ to arrive at the so-called Lee-Pomeransky (LP) representation Lee:2013hzt :
@@
We now show why Eq. (6) is equivalent to Eq. (2). The integral (6) can be regarded as the integration of a differential $n$-form on the $n$-dimensional hypersurface determined by $\mathcal{X}-\mathcal{Y}=0$:
@@
The differential $n$-form can be written as $f\,\omega$, where
@@
where the hat indicates that the corresponding factor is omitted, and
$f\equiv C(\bm{\nu}_{n})\left(\prod_{j=0}^{n}\alpha_{j}^{\nu_{j}-1}\right)\left(\alpha_{0}\,\mathcal{U}+\mathcal{F}\right)^{\lambda_{0}}\,.$ (10)
With the above definitions, the integral in Eq. (6) can be written as
$I(\bm{\nu}_{n})=\int_{S_{n}}f\,\omega\,.$ (11)
To show that the above relation is true, we use the fact that when restricted on $S_{n}$,
$\sum_{j=0}^{n}\mathrm{d}\alpha_{j}\frac{\partial(\mathcal{X}-\mathcal{Y})}{\partial\alpha_{j}}=\mathrm{d}(\mathcal{X}-\mathcal{Y})=0\,.$ (12)
Picking one of the variables (say, without loss of generality, $\alpha_{0}$) appearing in $\mathcal{X}-\mathcal{Y}$, we can express the integration measure $\omega$ on $S_{n}$ as
@@
This then leads to Eq. (11).
Note that $f$ is a homogeneous function of degree $(-n-1)$ in the variables $\{\alpha_{j}\}$. This can be used to demonstrate that the $n$-form $f\,\omega$ is closed, i.e., $\mathrm{d}(f\,\omega)=\mathrm{d}f\wedge\omega+f\,\mathrm{d}\omega=0$:
@@
Therefore, the integration of $\mathrm{d}(f\,\omega)$ in any region $\Delta_{n+1}$ in $\mathbb{R}^{n+1}$ is zero. By Stokes’ theorem, this means
@@
Now, suppose that $S_{n}$ and $S_{n}^{\prime}$ are two $n$-dimensional hypersurfaces defined by $\mathcal{X}-\mathcal{Y}=0$ and $\mathcal{X}^{\prime}-\mathcal{Y}^{\prime}=0$, respectively, and let $\Delta_{n+1}$ be the region enclosed by $S_{n}$, $S_{n}^{\prime}$ and the coordinate hyperplanes defined by $\alpha_{j}=0$ ($j=0,1,\ldots,n$). When restricted on the coordinate hyperplanes, $\omega$ always vanishes due to the factor of $\alpha_{j}$ in its definition (9). It is possible that we need to add hyperplanes at infinity to form a closed region. However, the integral at infinity does not contribute in dimensional regularization. Therefore, we can employ Eq. (15) to show that the integral is independent of the integration domain:
@@
Figure 1 illustrates the above idea for the case of two variables.
In the above, we have shown that the two contours $S_{1}$ and $S^{\prime}_{1}$ give rise to the same integral, and can be regarded as belonging to the same equivalence class. In the language of twisted homology (see, e.g., Aomoto:1414035 for more details), the equivalence classes of integration contours are elements (cycles) of a twisted homology group determined by the polynomial $\alpha_{0}\,\mathcal{U}+\mathcal{F}$. This homology group is dual to the twisted cohomology group of the integrands. In the literature Mizera:2017rqa ; Mastrolia:2018uzb , there have been extensive discussions on how to perform integral reduction using the vector-space structure of the cohomology groups. This can be done using the techniques of intersection theory zbMATH00713739 ; zbMATH01270294 ; zbMATH02112802 ; zbMATH06267077 ; zbMATH06447521 ; zbMATH06502598 ; zbMATH06454357 ; zbMATH07531013 ; zbMATH07527773 ; zbMATH07733418 . The homology groups of integration contours also admit a vector-space structure, and in principle can be used for integral reduction as well. However, this path has not been followed in the literature to the best of our knowledge.
@@ 2.3 The bubble family
We consider the reduction of $I(1,2)$. For that we introduce an auxiliary integral
@@
where $q_{1}\equiv(m_{1}^{2}+m_{2}^{2}-s)/(2m_{1}^{2})$. According to the result in Section 2.2, one can see that $g(1,1)$ is reducible to the sub-sector without $\alpha_{2}$. Applying the method from Section 2.2, we obtain
@@
Alternatively, we can make a variable change $\alpha_{1}\to\alpha_{1}+q_{1}\alpha_{2}$ and $\alpha_{2}\to\alpha_{2}(1-W\alpha_{2}/\alpha_{0})$ in Eq. (28), where $W\equiv\lambda(m_{1}^{2},m_{2}^{2},s)/\left[2(m_{1}^{2}-m_{2}^{2}+s)\right]$ and the Källén function is given by $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2yz-2zx$. We can then arrive at
@@
×$\quad\times\left[(\alpha_{1}+\alpha_{2})\alpha_{0}+\alpha_{1}\alpha_{2}(m_{1}^{2}+m_{2}^{2}-s)+\alpha_{1}^{2}m_{1}^{2}+\alpha_{2}^{2}m_{2}^{2}\right]^{\lambda_{0}}$
@@
×$\quad\times\left[(\alpha_{1}+\alpha_{2})\alpha_{0}+\alpha_{1}\alpha_{2}(m_{1}^{2}+m_{2}^{2}-s)+\alpha_{1}^{2}m_{1}^{2}+\alpha_{2}^{2}m_{2}^{2}\right]^{\lambda_{0}}\,.$ (31)
For the $h_{1}(1,1)$ function, we further apply the variable change $\alpha_{0}\to\alpha_{0}\alpha_{2}/(\alpha_{2}+\alpha_{1}/q_{1})$, $\alpha_{1}\to-\alpha_{1}\alpha_{2}/(\alpha_{2}+\alpha_{1}/q_{1})$, and denote $\alpha_{1}^{\prime}\equiv\alpha_{1}\left[1-W(2\alpha_{2}+\alpha_{1}/q_{1})/\alpha_{0}\right]$. We then obtain
@@ 3 The general method for one-loop integral reduction
The goal of one-loop integral reduction is to reduce the indices $\nu_{j}$ to either 00 or $1$. We now introduce an interesting technique to lower an index by one, at the cost of adding an auxiliary integration variable. While this seems to be meaningless at first sight, it will be employed in the derivation of the final reduction formula.
@@ 3.2 Reduction for irreducible sectors
In the following, we will aim for decreasing the power $\nu_{l}>1$ of the variable $\alpha_{l}$, where $l\in\{1,\cdots,n\}$. For that we will need to study the submatrix of $\bm{Z}$ obtained by removing the $l$-th row and $l$-th column. We will denote this submatrix as $\bm{Z}(\hat{l},\hat{l})$, and refer to it as “the $(l,l)$-submatrix”. A closely related concept is the $(i,j)$-minor of $\bm{Z}$, which is the determinant of the submatrix of $\bm{Z}$ without the $i$-th row and $j$-th column. The $(i,j)$-cofactor is further defined by multiplying the $(i,j)$-minor with $(-1)^{i+j}$. We will denote the $(i,j)$-cofactor of $\bm{Z}$ as $Z_{i,j}$.
The $(l,l)$-submatrix can be either non-singular with rank $n-1$, or singular with a smaller rank. In the following we will discuss the two situations separately.
@@ 3.2.1 When the $(l,l)$-submatrix is non-singular
where $\bm{\beta}^{\left(l\right)}$ denotes the vector obtained by replacing the $l$-th element (i.e., $\beta_{l}$) of the vector $\bm{\beta}\equiv(\beta_{1},\cdots,\beta_{n})^{T}$ with 00, and the elements of the vector $\bm{q}$ are given by
@@ 3.2.2 When the $(l,l)$-submatrix is singular
A special case occurs when the matrix $\bm{Z}$ has only one element, meaning that $\bm{Z}(\hat{l},\hat{l})$ is an empty matrix. In this case, we have $\nu=\nu_{1}$ and $\nu_{0}=d-\nu$. The procedure can stilled be carried out by defining $Z_{1,1}=1$, and we obtain the correct reduction formula for single-propagator integrals:
@@ 3.3.1 Reduction for general reducible sectors
×$\quad\times\int_{0}^{\infty}\mathrm{d}\alpha_{0}\,\alpha_{0}^{\nu_{0}-1}\left[\left(\mathcal{U}_{l,0}+\xi\alpha_{l}\right)\alpha_{0}+\mathcal{F}_{l,0}\right]^{\lambda_{0}}\,\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\,,$ (58)
@@ 3.3.2 Degenerate limits
where $\bm{x}_{n-1}$ denotes the sequence of $n-1$ variables, and $\bm{\eta}_{n-1}$ is the sequence of $n-1$ constants. Using the relation
@@ 3.4.1 Rewriting the integrals in shifted dimensions
where $\hat{\mathcal{O}}_{j}^{+}$ (for $\mathcal{O}\in\{\mathcal{U},\mathcal{F}\}$) represents an operator obtained by substituting each $\alpha_{k}$ in $\left(\partial\mathcal{O}/\partial\alpha_{j}\right)$ by $\hat{k}^{+}$. Note that Eq. (78) holds for general $L$-loop integrals, and for the purpose of this work, we can set $L=1\,$.
With Eq. (78), we can recursively increase a negative index to zero, at the cost of employing integrals in shifted dimensions. Let us consider a simple example: $I(-1,2)$ from the bubble integral family in Figure 3. Applying Eq. (78) with $j=1$, we obtain

### v2

@@ 1 Introduction
The paper is organized as follows. In Section 2, we introduce the equivalence relations of integral contours in Feynman parameterization and use several simple examples to demonstrate our reduction method. In Section 3, we present our general method for one-loop integral reduction, and provide recursive formulas that can be easily implemented in computer algebra. In Section 4, we demonstrate our method through several examples, including a preliminary extension to higher loops. In Section 5, we provide a summary and discuss the new challenges that may arise in future applications.
@@ 2.1 Feynman parametrization and equivalence classes of integration domains
where $\mathcal{X}$ and $\mathcal{Y}$ are non-negative homogeneous functions of degree 0 and $1$ in the variables $\alpha_{i}$}, respectively. We further require that, in the whole integration domain (i.e., $\mathcal{X}-\mathcal{Y}=0$ and all $\alpha_{j}\geq 0$), the inequality $\partial(\mathcal{X}-\mathcal{Y})/\partial\alpha_{j}\leq 0$ holds for every variable $\alpha_{j}$.
The original representation (2) then corresponds to the specific choice $\mathcal{X}=1$ and $\mathcal{Y}=\sum_{j\in\mathcal{S}}\alpha_{j}$. Note that by choosing $\mathcal{X}=1$ and $\mathcal{Y}=\alpha_{0}$, we can integrate out $\alpha_{0}$ to arrive at the so-called Lee-Pomeransky (LP) representation Lee:2013hzt :
@@
We now demonstrate that Eq. (6) is equivalent to Eq. (2) by showing that (6) is independent of the choices of $\mathcal{X}$ and $\mathcal{Y}$ (following an approach similar to that in Section 2.5.3 of Weinzierl:2022eaz ). The $\delta$-function in Eq. (6) effectively restricts the integration onto the $n$-dimensional hypersurface $S_{n}$ determined by $\mathcal{X}-\mathcal{Y}=0\,$:
@@
Let’s define an integrand function
$f\equiv C(\bm{\nu}_{n})\left(\prod_{j=0}^{n}\alpha_{j}^{\nu_{j}-1}\right)\left(\alpha_{0}\,\mathcal{U}+\mathcal{F}\right)^{\lambda_{0}}\,$, (9)
and an integration measure
@@
where the hat indicates that the corresponding factor is omitted. When restricted to the surface $S_{n}$, $\omega$ can be parametrized by $n$ out of $n+1$ variables. Picking one of the variables (say, without loss of generality, $\alpha_{0}$) appearing in $\mathcal{X}-\mathcal{Y}$, we can write
$\left.\omega\right|_{S_{n}}$ $=\left[\alpha_{0}\right]_{S_{n}}\mathrm{d}\alpha_{1}\wedge\cdots\wedge\mathrm{d}\alpha_{n}$
$+\sum_{j=1}^{n}\alpha_{j}\left[\left(\frac{\partial(\mathcal{X}-\mathcal{Y})}{\partial\alpha_{0}}\right)^{-1}\frac{\partial(\mathcal{X}-\mathcal{Y})}{\partial\alpha_{j}}\right]_{S_{n}}\mathrm{d}\alpha_{1}\wedge\ldots\wedge\mathrm{d}\alpha_{j}\wedge\ldots\wedge\mathrm{d}\alpha_{n}\,$, (11)
where the subscript $S_{n}$ means that $\alpha_{0}$ should be replaced by the solution to $\mathcal{X}-\mathcal{Y}=0$. Using the condition that $\partial(\mathcal{X}-\mathcal{Y})/\partial\alpha_{0}\leq 0$ when restricted to $S_{n}$, we arrive at
$\left.\omega\right|_{S_{n}}$ $=-\left|\frac{\partial(\mathcal{X}-\mathcal{Y})}{\partial\alpha_{0}}\right|^{-1}_{S_{n}}\left[\sum_{j=0}^{n}\alpha_{j}\frac{\partial(\mathcal{X}-\mathcal{Y})}{\partial\alpha_{j}}\right]_{S_{n}}\mathrm{d}\alpha_{1}\wedge\ldots\wedge\mathrm{d}\alpha_{n}\,.$ (12)
Next, we use the fact that $\mathcal{X}$ and $\mathcal{Y}$ are non-negative homogeneous functions of degree 0 and $1$, respectively. By Euler’s homogeneous function theorem, we have $\sum_{i}\alpha_{i}\partial_{\alpha_{i}}\mathcal{X}=0$ and $\sum_{i}\alpha_{i}\partial_{\alpha_{i}}\mathcal{Y}=\mathcal{Y}$, which imply
$\left[\sum_{j=0}^{n}\alpha_{j}\frac{\partial(\mathcal{X}-\mathcal{Y})}{\partial\alpha_{j}}\right]_{S_{n}}=\left.-\mathcal{Y}\right|_{S_{n}}=\left.-\mathcal{X}\right|_{S_{n}}\,.$ (13)
Combining the above information, we arrive at
@@
Hence, we see that the integral in Eq. (6) can be written as
$I(\bm{\nu}_{n})=\int_{S_{n}}f\,\omega\,.$ (15)
Suppose that $S_{n}$ and $S_{n}^{\prime}$ are two $n$-dimensional oriented hypersurfaces defined by $\mathcal{X}-\mathcal{Y}=0$ and $\mathcal{X}^{\prime}-\mathcal{Y}^{\prime}=0$, respectively. See Figure 1 for a simple illustration with $n=1$. We need to show that the integrals on $S_{n}$ and $S_{n}^{\prime}$ are the same. For that we will employ Stokes’ theorem. Let $\Delta_{n+1}$ be the region enclosed by $S_{n}$, $S_{n}^{\prime}$, the coordinate hyperplanes defined by $\alpha_{j}=0$ ($j=0,1,\ldots,n$), and possibly the hyperplanes at infinity. Note that $f$ is a homogeneous function of degree (n1)(-n-1) in the variables $\{\alpha_{j}\}$. This can be used to demonstrate that the $n$-form $f\,\omega$ is closed, i.e., $\mathrm{d}(f\,\omega)=\mathrm{d}f\wedge\omega+f\,\mathrm{d}\omega=0$:
@@
Therefore, the integration of $\mathrm{d}(f\,\omega)$ in any region $\Delta_{n+1}$ in $\mathbb{R}^{n+1}$ is zero. By Stokes’ theorem, this means that
@@
When restricted on the coordinate hyperplanes as part of $\partial\Delta_{n+1}$, $\omega$ always vanishes due to the factor of $\alpha_{j}$ in its definition (10). For the hyperplanes at infinity, we recall that the inequality $\partial(\mathcal{X}-\mathcal{Y})/\partial\alpha_{j}]_{S_{n}}\leq 0$ holds for every variable $\alpha_{j}\,$. Therefore, if the integrand has no singularities on the hyperplanes at infinity, its integral is confined to an infinitesimally small solid angle and thus yields a vanishing contribution. If, on the other hand, the integrand exhibits singularities on the hyperplanes at infinity, one needs to introduce regulators to define the integral. In dimensional regularization, the integral is regarded as an analytic function of the complex variable $\lambda_{0}=-d/2\,$. Hence, when some $\alpha_{j}\to\infty\,$, there always exists some region of $\lambda_{0}$ where $\alpha_{0}^{-(L+1)\lambda_{0}}(\alpha_{0}\,\mathcal{U}+\mathcal{F})^{\lambda_{0}}$ goes to zero sufficiently fast. It is then clear that the integrals on hyperplanes at infinity vanish for any $\lambda_{0}$ (including integer dimensions) by analytic continuation. The possible introduction of analytic regulators $\alpha_{j}^{\rho}$ does not change this conclusion, since $\rho$ should then be regarded as another complex variable on which the integral depends. Putting all the above together, we are now left with only $S_{n}$ and $S_{n}^{\prime}$ in $\partial\Delta_{n+1}\,$. After taking care of the orientations of the hypersurfaces, we finally arrive at
@@
In the above, we have shown that the two contours $S_{n}$ and $S^{\prime}_{n}$ give rise to the same integral, and can be regarded as belonging to the same equivalence class. In the language of twisted homology (see, e.g., Aomoto:1414035 for more details), the equivalence classes of integration contours are elements (cycles) of a twisted homology group determined by the polynomial $\alpha_{0}\,\mathcal{U}+\mathcal{F}$. This homology group is dual to the twisted cohomology group of the integrands. In the literature Mizera:2017rqa ; Mastrolia:2018uzb , there have been extensive discussions on how to perform integral reduction using the vector-space structure of the cohomology groups. This can be done using the techniques of intersection theory zbMATH00713739 ; zbMATH01270294 ; zbMATH02112802 ; zbMATH06267077 ; zbMATH06447521 ; zbMATH06502598 ; zbMATH06454357 ; zbMATH07531013 ; zbMATH07527773 ; zbMATH07733418 . The homology groups of integration contours also admit a vector-space structure, and in principle can be used for integral reduction as well. However, this path has not been followed in the literature to the best of our knowledge.
@@ 2.3 The bubble family
We consider the reduction of $I(1,2)$. Note that the integral shown in Eq. (29) for positive $\nu_{1}$ and $\nu_{2}$ does not possess the property of being reducible to sub-sectors as described in Section 2.2. Therefore, we need to split the integral into two parts: one part that is reducible to sub-sectors, and the other part that can be eventually related to $I(1,1)$. To this end, we consider the auxiliary integral
@@
where $q_{1}\equiv(m_{1}^{2}+m_{2}^{2}-s)/(2m_{1}^{2})$. The motivation for considering the above integral comes from two perspectives. First of all, the integral itself is reducible. According to the result in Section 2.2, one can see that $g(1,1)$ is reducible to the sub-sector without $\alpha_{2}$. Applying the method from Section 2.2, we obtain
@@
On the other hand, the integral can be related to a linear combination of $I(1,1)$ and $I(1,2)$ through a change of variables. We perform the variable change $\alpha_{1}\to\alpha_{1}+q_{1}\alpha_{2}$ and $\alpha_{2}\to\alpha_{2}(1-W\alpha_{2}/\alpha_{0})$ in Eq. (30), where $W\equiv\lambda(m_{1}^{2},m_{2}^{2},s)/\left[2(m_{1}^{2}-m_{2}^{2}+s)\right]$ and the Källén function is given by $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2yz-2zx$. We can then arrive at
@@
$\quad\times\left[(\alpha_{1}+\alpha_{2})\alpha_{0}+\alpha_{1}\alpha_{2}(m_{1}^{2}+m_{2}^{2}-s)+\alpha_{1}^{2}m_{1}^{2}+\alpha_{2}^{2}m_{2}^{2}\right]^{\lambda_{0}}$
@@
$\quad\times\left[(\alpha_{1}+\alpha_{2})\alpha_{0}+\alpha_{1}\alpha_{2}(m_{1}^{2}+m_{2}^{2}-s)+\alpha_{1}^{2}m_{1}^{2}+\alpha_{2}^{2}m_{2}^{2}\right]^{\lambda_{0}}\,.$ (33)
The $h_{1}(1,1)$ function is also reducible to sub-sectors. This can be seen by further applying the variable change $\alpha_{0}\to\alpha_{0}\alpha_{2}/(\alpha_{2}+\alpha_{1}/q_{1})$, $\alpha_{1}\to-\alpha_{1}\alpha_{2}/(\alpha_{2}+\alpha_{1}/q_{1})$. Denoting $\alpha_{1}^{\prime}\equiv\alpha_{1}\left[1-W(2\alpha_{2}+\alpha_{1}/q_{1})/\alpha_{0}\right]$, we have
@@ 3 The general method for one-loop integral reduction
The goal of one-loop integral reduction is to reduce the indices $\nu_{j}$ to either 0 or $1$. We now introduce an interesting technique to lower an index by one, at the cost of adding an auxiliary integration variable. While this seems to be meaningless at first sight, it will be employed in the derivation of the final reduction formula.
@@ 3.2 Reduction for irreducible sectors
[…]
The $(l,l)$-submatrix can be either non-singular with rank n1n-1, or singular with a smaller rank. In the following we will discuss the two situations separately.
@@ 3.2.1 When the $(l,l)$-submatrix is non-singular
where $\bm{\beta}^{\left(l\right)}$ denotes the vector obtained by replacing the $l$-th element (i.e., $\beta_{l}$) of the vector $\bm{\beta}\equiv(\beta_{1},\cdots,\beta_{n})^{T}$ with 0, and the elements of the vector $\bm{q}$ are given by
@@ 3.2.2 When the $(l,l)$-submatrix is singular
[…]
@@ 3.3.1 Reduction for general reducible sectors
$\quad\times\int_{0}^{\infty}\mathrm{d}\alpha_{0}\,\alpha_{0}^{\nu_{0}-1}\left[\left(\mathcal{U}_{l,0}+\xi\alpha_{l}\right)\alpha_{0}+\mathcal{F}_{l,0}\right]^{\lambda_{0}}\,\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\,$, (60)
@@ 3.3.2 Degenerate limits
where $\bm{x}_{n-1}$ denotes the sequence of n1n-1 variables, and $\bm{\eta}_{n-1}$ is the sequence of n1n-1 constants. Using the relation
@@ 3.4.1 Rewriting the integrals in shifted dimensions
where $\hat{\mathcal{O}}_{j}^{+}$ (for $\mathcal{O}\in\{\mathcal{U},\mathcal{F}\}$) represents an operator obtained by substituting each $\alpha_{k}$ in $\left(\partial\mathcal{O}/\partial\alpha_{j}\right)$ by $\hat{k}^{+}$. Note that Eq. (80) holds for general $L$-loop integrals, and for the purpose of this work, we can set L=1L=1\,.
With Eq. (80), we can recursively increase a negative index to zero, at the cost of employing integrals in shifted dimensions. Let us consider a simple example: I(1,2)I(-1,2) from the bubble integral family in Figure 3. Applying Eq. (80) with $j=1$, we obtain